Find the total surface area of a tent shaped object as follows:
Base is 12 X 6
Each rectangle is 12 X 8.54
Each triangle is 6 X 8.54 with a height of 8

Answer by checkley77(12844)

12*6=72 ft^2
2(12*8.54)=2*102.48=204.96
2(6*8)/2)=48
72+102.48+48=224,48 ft^2.